# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ 2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
The proof strategy of Theorem 2.2 is to leverage the Poisson equation ${\bm{G}}-\left({{\mathrm{Q}}{\bm{G}}}\right)={\bm{F}}$ to construct a matrix martingale ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{G}}(Z_{j})-({\mathrm{Q}}{\bm{G}})(Z_{j-1})$, and then we apply our Rosenthal-Burkholder inequality (Theorem 2.1). This method is first introduced by Durmus et al. (2023), who derived a Hilbert space-valued Rosenthal inequality for Markov chains. A detailed outline of the proof can be found in Section 4.2.
For the dimension-free results, as $n\to\infty$, we can show that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|)$, which is a natural dimension factor in this problem.
@@
Fortunately, if we do not insist on obtaining the result with a leading term matching the Markov chain CLT, we can still derive results with the sharp dependence on $t_{\operatorname{mix}}$ by using the blocking argument. In the following theorem, we present another version of the Rosenthal inequality, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}^{2}$, which is an upper bound for the long-run variance $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$.
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{16e}{3}C_{R,1}\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+\left({\frac{16e}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n}.$
@@
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $\frac{16e}{3}C_{R,1}\sqrt{\frac{\log(\delta^{-1}{\vee}d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+D_{3}\log(\delta^{-1}{\vee}d)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n},$
where $D_{3}>0$ is a universal constant, which can be taken as $6eC_{R,2}+eD_{2}$.
@@
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $eC_{R,1}\sqrt{\frac{\log(\delta^{-1}{\vee}d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+e(D_{1}+D_{2})t_{\operatorname{mix}}^{1/2}\left({\frac{\log(\delta^{-1}{\vee}d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
Discussion.
Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) derived matrix concentration inequalities for Markov chains with a non-zero absolute spectral gap. Instead of the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004) used in these works, we leverage the Poisson equation approach (Durmus et al., 2023), which allows us to overcome several limitations of the previous results: First, our results can be applied to general non-reversible Markov chains without assuming a non-zero absolute spectral gap. Second, our results can be extended to more general ergodicity assumptions, under which we do not need to assume the boundedness of functions, as detailed in Section 2.2.1. Third, the leading term of our results matches the Markov chain CLT, rather than relying on a variance proxy, e.g. $t_{\operatorname{mix}}\|{\mathbb{E}}_{\pi}[{\bm{F}}^{2}(Z_{0})]\|$ or $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|^{2}$. Finally, under further assumptions, we obtain dimension-free results, which the previous works, which used the multi-matrix Golden-Thompson inequality, fail to achieve them. Applying the dimension-free results, we can also obtain the concentration inequalities for linear operators in infinite-dimensional Hilbert spaces.
Our results have the following limitations: As mentioned before, the residual term in the results where the leading term matches the Markov chain CLT (Theorem 2.2, Theorem 2.2, Theorem 2.4) has a suboptimal dependence on $t_{\operatorname{mix}}$, failing to achieve the sharp dependence on $t_{\operatorname{mix}}$ in the residual term as in the Hilbert space-valued version of the inequality (Durmus et al., 2023). We leave the improvement to the sharp dependence on $t_{\operatorname{mix}}$ as future work. Another limitation is that the Poisson equation approach is only applicable to time-homogeneous functions (additive functionals of Markov chains), whereas the results in Neeman et al. (2024) apply to time-inhomogeneous functions. This makes our results inapplicable to problems such as stochastic approximation with Markovian data (Li et al., 2023; Mou et al., 2024; Huo et al., 2024; Samsonov et al., 2025). We also leave the extension to time-inhomogeneous functions as future work.
2.2.1 Extension: Results for Geometrically $V$-Ergodic Markov Chains
We now introduce the more general geometrically $V$-ergodic Markov chains (Meyn and Tweedie, 2012; Douc et al., 2018; Durmus et al., 2023) and present matrix Rosenthal inequalities under the ergodicity assumption. Let $V\colon{\mathcal{Z}}\to[e,\infty)$ be a Lyapunov function satisfying $\pi(V)<\infty$. We denote by ${\mathcal{P}}_{V}$ the set of all probability distributions whose elements $\xi$ satisfy $\xi(V)<\infty$. For any real-valued measurable function $f$, we also define its $V$-norm as $\|f\|_{V}=\sup_{z\in{\mathcal{Z}}}{\left|{f(z)}\right|}/{V(z)}$. We will see that, under the ergodicity assumption, for matrix functions $\bm{F}$} whose growth rate does not exceed $V^{1/p}$, the convergence of $n^{-1}{\bm{S}}_{n}$ in the $L^{p}$ norm can be obtained, which allows us to extend the result to unbounded functions. In this section we omit the results in the rectangular case.
@@ 2.2.1 Extension: Results for Geometrically $V$-Ergodic Markov Chains
Assumption 2.2 is equivalent to the statement that for any $\xi\in{\mathcal{P}}_{V}$ and $k\in{\mathbb{N}}$,
$\left\|{\xi{\mathrm{Q}}^{k}-\pi}\right\|_{V}\leq\varkappa\left({\xi(V)+\pi(V)}\right)\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}.$
Hence, Assumption 2.1 is a special case of Assumption 2.2 with $V\equiv e$ and $\varkappa=1$. For more details on the total variation norm and the $V$-total variation norm, one can refer to Douc et al. (2018, Appendix D).
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+D_{4}p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{-3/4}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+D_{5}p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{-(1-1/p)}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\pi(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
where $C_{R,1}$ is defined in Theorem 2.1 and $D_{4},D_{5}>0$ are universal constants, which can be traced in the proof.
@@
Next, we present a parallel result to Theorem 2.3, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}^{2}$.
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{8}{3}C_{R,1}p\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{11}{3}}\right)\frac{pt_{\operatorname{mix}}}{n^{1-1/p}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
The proof can be found in Lemma 4.3 (b).
We can also use Markov inequality to obtain high-probability upper bounds similar to Corollary 2.1 and Corollary $2.$ For brevity, we omit them here.
@@ 3 Applications
We use the same notation as in Section 2.2. Suppose $(Z_{k})_{k\in{\mathbb{N}}}$ is a uniformly geometrically ergodic stationary Markov chain (Assumption 2.1) and $\bm{f}\colon{\mathcal{Z}}\to{\mathbb{R}}^{d}$ is a bounded vector-valued function. We first consider the problem of covariance estimation on Markovian data, which is also considered in Fan et al. (2021). That is, we are interested in analyzing the spectral norm error in estimating the covariance matrix ${\bm{\Sigma}}=\pi(\bm{f}\bm{f}^{\top})={\mathbb{E}}_{\pi}[\bm{f}(Z_{0})\bm{f}(Z_{0})^{\top}]$ using the sample covariance matrix $\hat{{\bm{\Sigma}}}_{n}=n^{-1}\sum_{k=0}^{n-1}\bm{f}(Z_{k})\bm{f}(Z_{k})^{\top}$. Let ${\bm{F}}(z):=\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}}$. Then $\bm{F}$} satisfies $\pi({\bm{F}})=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}\leq\|\|\bm{f}\|\|_{\infty}^{2}$. Using the matrix Bernstein inequality for Markov chains (Corollary 2.2), we obtain that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\frac{1}{n}\sup_{k\leq n}\left\|{\sum_{j=0}^{k-1}\left({\bm{f}(Z_{j})\bm{f}(Z_{j})^{\top}-{\bm{\Sigma}}}\right)}\right\|\leq C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right),$ (9)
where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance of $\bm{F}$} defined in Eqn. (2).
@@
$1-\left({{\bm{v}}_{1}^{\top}\hat{{\bm{v}}}_{1,n}}\right)^{2}\leq C\frac{\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{(\lambda_{1}-\lambda_{2})^{2}n}+O\left({\frac{1}{n^{3/2}}}\right),$
@@
Under further assumptions on $\bm{f}$, we can obtain dimension-free results. Let ${\bm{\Upsilon}}\in{\mathbb{H}}_{d}^{+}$ be a real PSD matrix with spectral decomposition ${\bm{\Upsilon}}=\sum_{i=1}^{d}\mu_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{\top}$ such that $\mu_{1}\geq\mu_{2}\geq\ldots\geq\mu_{d}\geq 0$, and $\bm{f}(z)=\sum_{i=1}^{d}\alpha_{i}(z){\bm{u}}_{i}$—the representation of $\bm{f}$ on the orthonormal basis $\{{\bm{u}}_{1},\ldots,{\bm{u}}_{d}\}$. Suppose that $\sup_{z\in{\mathcal{Z}}}\sum_{i=1}^{d}\alpha_{i}^{2}(z)/\mu_{i}\leq 1$ ($0/0:=0$ and $1/0:=\infty$). Then by the Schur complement lemma (Zhang, 2006), one has that $\bm{f}(z)\bm{f}(z)^{\top}\preccurlyeq{\bm{\Upsilon}}$ for any $z\in{\mathcal{Z}}$. Furthermore, ${\bm{F}}^{2}(z)=(\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}})^{2}\preccurlyeq 2[\bm{f}(z)\bm{f}(z)^{\top}]^{2}+2{\bm{\Sigma}}^{2}\preccurlyeq 2\|\|\bm{f}\|\|_{\infty}^{2}[\bm{f}(z)\bm{f}(z)^{\top}+{\bm{\Sigma}}]\preccurlyeq 4\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}$. Using the dimension-free results, we can replace the ambient $d$ in the bounds by the effective rank $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,(1/2)\log(1/\delta)}])$. Similar results hold for the more general geometrically $V$-ergodic Markov chains and unbounded function $\bm{f}$, and we omit them for brevity.
Compared with the results in Kumar and Sarkar (2023); Neeman et al. (2024), our results do not require the Markov chain to have an absolute spectral gap, while obtaining a leading term that matches the Markov chain CLT and gaining dimension-free results. Moreover, our results can be extended to more general ergodicity assumptions and unbounded functions. These aspects demonstrate the advantages of our results over the previous ones as we discussed earlier.

### v2

@@ 2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
The proof strategy of Theorem 2.3 is to leverage the Poisson equation ${\bm{G}}-\left({{\mathrm{Q}}{\bm{G}}}\right)={\bm{F}}$ to construct a matrix martingale ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{G}}(Z_{j})-({\mathrm{Q}}{\bm{G}})(Z_{j-1})$, and then we apply our Rosenthal-Burkholder inequality (Theorem 2.1). This method was introduced by Durmus et al. (2023), who derived a Hilbert space-valued Rosenthal inequality for Markov chains. A detailed outline of the proof can be found in Section 4.2.
@@
Fortunately, if we do not insist on obtaining the result with a leading term matching the Markov chain CLT, we can still derive results with the sharp dependence on $t_{\operatorname{mix}}$ by using the blocking argument. In the following theorem, we present another version of the Rosenthal inequality, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{\infty}^{2}$, which is an upper bound for the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$.
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{16e}{3}C_{1}\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+\left({\frac{16e}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)\frac{t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n}.$
@@
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $\frac{16e}{3}C_{1}\sqrt{\frac{\log(d/\delta)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+D_{3}\frac{\log(d/\delta)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}}{n},$
where $D_{3}>0$ is a universal constant, which can be taken as $6eC_{2}+eD_{2}$.
@@
$\frac{1}{n}\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|\leq$ $eC_{1}\sqrt{\frac{\log(d/\delta)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+e(D_{1}+D_{2})t_{\operatorname{mix}}^{1/2}\left({\frac{\log(d/\delta)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
2.2.1 Results for Geometrically $V$-Ergodic Markov Chains
The assumptions of uniform geometric ergodicity (Assumption 2.1) and bounded functions are excessively strong and fail to hold in many applications. We now introduce the more general geometrically ergodic Markov chains with Lyapunov function $V$ and present matrix Rosenthal inequalities under the ergodicity assumption. Let $V\colon{\mathcal{Z}}\to[e,\infty)$ be a Lyapunov function satisfying $\mu(V)<\infty$. We denote by ${\mathcal{P}}_{V}$ the set of all probability distributions whose elements $\xi$ satisfy $\xi(V)<\infty$. For any real-valued measurable function $f$, we also define its $V$-norm as $\|f\|_{V}=\sup_{z\in{\mathcal{Z}}}{\left|{f(z)}\right|}/{V(z)}$. We will see that, under the ergodicity assumption, for matrix functions $\bm{F}$} whose growth rate does not exceed $V^{1/p}$, the convergence of $n^{-1}{\bm{S}}_{n}$ in the $L^{p}$ norm can be obtained, which allows us to extend the result to unbounded functions. In this section we omit the results in the rectangular case for brevity.
@@ 2.2.1 Results for Geometrically $V$-Ergodic Markov Chains
This assumption can be verified via the Foster-Lyapunov drift condition (Meyn and Tweedie, 2012). Assumption 2.2 implies that for any $\xi\in{\mathcal{P}}_{V}$ and $k\in{\mathbb{N}}$,
$\left\|{\xi{\mathrm{Q}}^{k}-\mu}\right\|_{V}\leq\varkappa\left({\xi(V)+\mu(V)}\right)\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor}.$
Hence, Assumption 2.1 is a special case of Assumption 2.2 with $V\equiv e$ and $\varkappa=1$.
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+D_{4}p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{-3/4}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+D_{5}p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{-(1-1/p)}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\mu(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}},$
where $C_{1}$ is defined in Theorem 2.1 and $D_{4},D_{5}>0$ are universal constants, which can be traced in the proof.
@@
Next, we present a parallel result to Theorem 2.4, whose leading term is given by the variance proxy $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}^{2}$, and the proof can be found in Lemma 4.3 (b).
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{8}{3}C_{1}p\sqrt{\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{11}{3}}\right)\frac{pt_{\operatorname{mix}}}{n^{1-1/p}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
We can also use Markov inequality to obtain high-probability upper bounds similar to Corollary 2.1 and Corollary 2.2. For brevity, we omit them here.
Discussion.
As discussed in Related Work, Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) derived matrix concentration inequalities for Markov chains with a non-zero absolute $L^{2}$-spectral gap; and Van Werde and Sanders (2025) derived universality-based concentration inequalities for $\psi$-mixing Markov chains. Instead of the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018), the León-Perron argument (León and Perron, 2004), or the universality-based approach (Bandeira et al., 2023; Brailovskaya and van Handel, 2024) used in these works, we leverage the Poisson equation approach (Durmus et al., 2023), which allows us to overcome several limitations of the previous results: First, our results can be applied to general non-reversible geometrically ergodic Markov chains without assuming a non-zero absolute spectral gap or $\psi$-mixing condition. Second, we do not need to assume the boundedness of functions under the geometric ergodicity assumption. Third, the leading term of our results matches the Markov chain CLT, rather than relying on a sub-optimal variance proxy as in (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025). Finally, under further assumptions, we obtain dimension-free results, while the approaches employed in previous works failed to achieve them. Applying the dimension-free results, we can also obtain the concentration inequalities for linear operators in infinite-dimensional Hilbert spaces.
Nevertheless, our results have the following limitations: As mentioned before, the residual term in the results where the leading term matches the Markov chain CLT (Theorem 2.3, Theorem 2.2, Theorem 2.5) has a suboptimal dependence on $t_{\operatorname{mix}}$, failing to achieve the sharp dependence on $t_{\operatorname{mix}}$ in the residual term. This issue arises because operator convexity (monotonicity) for a scalar-valued function acted on matrices imposes stronger requirements than convexity (monotonicity), and consequently, arguments from the scalar case (Durmus et al., 2023) fail to hold. We leave the improvement to the sharp dependence on $t_{\operatorname{mix}}$ for future work. Another limitation is that the Poisson equation approach is only applicable to time-homogeneous functions (additive functionals of Markov chains), whereas the results in Neeman et al. (2024) apply to time-inhomogeneous functions. This makes our results inapplicable to problems such as stochastic approximation with Markovian data (Li et al., 2023; Mou et al., 2024; Huo et al., 2024; Samsonov et al., 2025). We also leave the extension to time-inhomogeneous functions as future work. A final limitation is that the universality-based inequalities in Van Werde and Sanders (2025) yield sharper bounds, although they require the stronger $\psi$-mixing condition and a finite dimension $d$. A key direction for future research is to determine if universality-based sharp concentration inequalities can be derived under weaker ergodicity assumptions. One potential approach is to first establish a universality-based moment inequality for matrix martingales and then apply the Poisson equation method. We leave this as future work as well.
@@ 3 Applications
We use the same notation as in Section 2.2. Suppose $(Z_{k})_{k\in{\mathbb{N}}}$ is a uniformly geometrically ergodic stationary Markov chain (Assumption 2.1) and $\bm{f}\colon{\mathcal{Z}}\to{\mathbb{R}}^{d}$ is a bounded vector-valued function. We first consider the problem of covariance estimation on Markovian data, which is also considered in Fan et al. (2021). That is, we are interested in analyzing the spectral norm error in estimating the covariance matrix ${\bm{\Sigma}}=\mu(\bm{f}\bm{f}^{\top})={\mathbb{E}}_{\mu}[\bm{f}(Z_{0})\bm{f}(Z_{0})^{\top}]$ using the sample covariance matrix $\hat{{\bm{\Sigma}}}_{n}=n^{-1}\sum_{k=0}^{n-1}\bm{f}(Z_{k})\bm{f}(Z_{k})^{\top}$. Let ${\bm{F}}(z):=\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}}$. Then $\bm{F}$} satisfies $\mu({\bm{F}})=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}\leq\|\|\bm{f}\|\|_{\infty}^{2}$. Using the matrix Bernstein inequality for Markov chains (Corollary 2.2), we obtain that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
[…]
where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance of $\bm{F}$} defined in Eqn. (3).
@@
$1-\left({{\bm{v}}_{1}^{\top}\hat{{\bm{v}}}_{1,n}}\right)^{2}\leq C\frac{\log(d/\delta)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{(\lambda_{1}-\lambda_{2})^{2}n}+O\left({\frac{1}{n^{3/2}}}\right),$
@@
Under further assumptions on $\bm{f}$, we can obtain dimension-free results. Let ${\bm{\Upsilon}}\in{\mathbb{H}}_{d}^{+}$ be a real PSD matrix with spectral decomposition ${\bm{\Upsilon}}=\sum_{i=1}^{d}\alpha_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{\top}$ such that $\alpha_{1}\geq\alpha_{2}\geq\ldots\geq\alpha_{d}\geq 0$, and $\bm{f}(z)=\sum_{i=1}^{d}\beta_{i}(z){\bm{u}}_{i}$—the representation of $\bm{f}$ on the orthonormal basis $\{{\bm{u}}_{1},\ldots,{\bm{u}}_{d}\}$. Suppose that $\sup_{z\in{\mathcal{Z}}}\sum_{i=1}^{d}\beta_{i}^{2}(z)/\alpha_{i}\leq 1$ ($0/0:=0$ and $1/0:=\infty$). Then by the Schur complement lemma (Zhang, 2006), one has that $\bm{f}(z)\bm{f}(z)^{\top}\preccurlyeq{\bm{\Upsilon}}$ for any $z\in{\mathcal{Z}}$. Furthermore, ${\bm{F}}^{2}(z)=(\bm{f}(z)\bm{f}(z)^{\top}-{\bm{\Sigma}})^{2}\preccurlyeq 2[\bm{f}(z)\bm{f}(z)^{\top}]^{2}+2{\bm{\Sigma}}^{2}\preccurlyeq 2\|\|\bm{f}\|\|_{\infty}^{2}[\bm{f}(z)\bm{f}(z)^{\top}+{\bm{\Sigma}}]\preccurlyeq 4\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}$. Using the dimension-free results, we can replace the ambient $d$ in the bounds by the effective rank $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}\|\|\bm{f}\|\|_{\infty}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,(1/2)\log(1/\delta)}])$. Similar results hold for the more general geometrically $V$-ergodic Markov chains and unbounded function $\bm{f}$, and we omit them for brevity.
Compared with the results in Kumar and Sarkar (2023); Neeman et al. (2024), our results obtain a leading term that matches the Markov chain CLT and gain dimension-free results. Moreover, our results can be extended to the more general geometrically $V$-ergodicity assumption and unbounded functions. These aspects demonstrate the advantages of our results over the previous ones as we discussed earlier.